### v1

# WarriorCoder: Learning from Expert Battles to Augment Code Large Language Models

Abstract
Despite recent progress achieved by code large language models (LLMs), their remarkable abilities are largely dependent on fine-tuning on the high-quality data, posing challenges for data collection and annotation. To address this, current methods often design various data flywheels to gather complex code instructions, enabling models to handle more intricate tasks. However, these approaches typically rely on off-the-shelf datasets and data augmentation from the limited pool of proprietary LLMs (e.g., Claude, GPT4, and so on), which limits the diversity of the constructed data and makes it prone to systemic biases. In this paper, we propose WarriorCoder which learns from expert battles to address these limitations. Specifically, we create an arena for current expert code LLMs, where each model challenges and responds to others’ challenges, with evaluations conducted by uninvolved judge models. This competitive framework generates novel training data constructed from scratch, harnessing the strengths of all participants. Experimental results demonstrate that WarriorCoder achieves competitive performance compared to previous methods, even without relying on proprietary LLMs.
WarriorCoder: Learning from Expert Battles to Augment Code Large Language Models
Huawen Feng1,2, Pu Zhao2, Qingfeng Sun2, Can Xu2, Fangkai Yang2, Lu Wang2, Qianli Ma1, Qingwei Lin2, Saravan Rajmohan2, Dongmei Zhang2, Qi Zhang2 1School of Computer Science and Engineering, South China University of Technology, China 2Microsoft qianlima@scut.edu.cn {v-huawenfeng,puzhao,qins,caxu,fangkai.yang,wlu,qlin,saravan.rajmohan, dongmeiz,zhang.qi}@microsoft.com
1 Introduction
Recent large language models (LLMs) have demonstrated impressive performance on code-related tasks Li et al. (2023); Rozière et al. (2023); Guo et al. (2024); DeepSeek-AI et al. (2024); Li et al. (2022); Nijkamp et al. (2023); Zheng et al. (2023b); Fried et al. (2023); Wang et al. (2021). These successes highlight that pre-training on vast amounts of code data significantly enhances their core coding abilities. In addition to pre-training, several approaches that fine-tune LLMs with instruction-following data Ding et al. (2023) have also made substantial progress in improving models’ understanding of user instructions and the quality of their responses. However, the effectiveness of post-training is heavily dependent on the availability of high-quality data Xu et al. (2024a), and challenges of data collection and annotation remain difficult to overcome.
Figure 1: The comparisons between our method and traditional data flywheels. Different from the previous work, we create an arena for current expert code LLMs. No demand for seed datasets, human prompts and proprietary models’ annotations, the target model learns from the pairwise competitions and integrates the strengths of competitors.
To address these challenges, some approaches propose various data flywheels to generate instruction data. Building on Self-Instruct, Chaudhary (2023) constructs Code Alpaca by prompting teacher LLMs to generate instructions in a few-shot setting. To further enhance the diversity and complexity of Code Alpaca, WizardCoder Luo et al. (2024b) employs Evol-Instruct to evolve the original instructions. These methods apply general data augmentation to instruction construction, lacking specific design considerations for the code domain. Given that, recent methods specifically design frameworks for instruction generation tailored to code. For example, WaveCoder Yu et al. (2024) collects raw code snippets and defines different tasks based on them. Similarly, InverseCoder Wu et al. (2024b) prompts code LLMs to generate high-quality instructions based on the original code through techniques like code summarization and self-evaluation. These methods construct data for code in various ways, effectively enhancing the model’s code generation capabilities. However, they still rely on existing datasets Muennighoff et al. (2024) and calls to proprietary LLMs (e.g., GPT-3.5, GPT-4, etc.), making data collection costly. Additionally, the limited data sources and annotators constrain the diversity Yu et al. (2023); Wang et al. (2023) of the data and inherit the system biases inherent in the limited pool of annotators Wei et al. (2024).
The challenges mentioned above have motivated us to propose WarriorCoder which learns from expert battles to address the current limitations. Specifically, we create an arena for current expert code LLMs and organize competitions among them. As shown in Figure 1, the models compete against each other, with evaluations conducted by uninvolved models. The attacker challenges all other models in its area of expertise, and the target model learns from the winner of these pairwise competitions. In this way, we can generate novel training data that incorporates the strengths of all the expert LLMs. Different from the previous work, our training data is fully automatically constructed from scratch, rather than expanding available datasets by prompting limited proprietary LLMs. In essence, the target model absorbs the best practices from a variety of powerful code LLMs, enabling it to be theoretically extended indefinitely to include emerging expert code LLMs. Moreover, our method eliminates the need for human involvement and proprietary LLMs in the training process, making it possible to collect high-quality, diverse data at a low cost. The main contributions of this paper are summarized as follows:
  • We identify the limitations of current data flywheels through detailed analysis and propose WarriorCoder, a new paradigm in which the target model absorbs best practices from a variety of strong code LLMs, rather than learning from existing datasets or their expansions.
  • We create an arena for current strong LLMs and generate novel training data that integrate the strengths of all participants, which can be easily scaled to accommodate the growth of expert code LLMs.
  • Extensive experiments demonstrate the effectiveness of WarriorCoder, with ablation and comparison studies explaining how and why it works. It achieves competitive performance compared to previous methods, even without relying on proprietary LLMs.
2 Related Work
2.1 Code LLMs
Code plays a crucial role in application areas for LLMs Jain et al. (2024), attracting significant interest from both academia and industry. Codex Chen et al. (2021), an LLM with 12 billion parameters, can solve 72.31% of complex Python programming problems posed by humans. Following the success of Codex Lyu et al. (2024), the rise of new code LLMs has demonstrated even greater capabilities, such as code generation and debugging, as model sizes continue to grow Hou et al. (2023). These models have become indispensable in a variety of code-specific tasks Zan et al. (2023). Despite this impressive progress, the performance of current open-source models still lags behind that of proprietary ones (e.g., GPT-3.5, GPT-4, etc.), primarily because stronger models often keep their training data proprietary Hui et al. (2024). As a result, the lack of publicly available code datasets remains a significant barrier to further development in this field.
2.2 Learning from Battles
Studying how people interact with LLMs in real-world scenarios is a pressing need for ensuring the alignment of LLMs Chiang et al. (2024). The LMSYS Chatbot Arena Zheng et al. (2024) has emerged as a groundbreaking initiative for exploring real-world LLM-user interactions, collecting and analyzing data from an open platform with over 240K votes. Experimental results demonstrate that the quality of data from the LMSYS Chatbot Arena is competitive with that of ShareGPT Chiang et al. (2023), underscoring its value for training. Now more and more attentions are paid on learning from the battles between LLMs Li et al. (2024); Myrzakhan et al. (2024); Bogomolov et al. (2024). However, collecting data through human online evaluations is both expensive and time-consuming. To address this, recent work has leveraged LLMs to provide their preferences when faced with different responses Luo et al. (2024a); Zhao et al. (2024). Although these methods eliminate the need for human annotation during data collection, they still require pre-designed, high-quality instructions.
2.3 LLM as a Judge
Offering an automatic alternative to the scalability challenges inherent in human evaluation, the concept of LLM-as-a-judge has garnered significant public attention in recent years Chiang and Lee (2023). As large language models (LLMs) such as GPT-4 have demonstrated impressive capabilities, they are increasingly being considered for use in evaluating other machine-generated outputs Weyssow et al. (2024). Experimental results by Thakur et al. (2024) show that these strong LLMs are capable of achieving a Cohen’s Kappa coefficient over 80%, a metric typically used to assess the level of agreement between human raters. This performance level is comparable to the consensus found among human experts, highlighting the potential of LLMs to serve as reliable evaluators in various contexts. However, judge models often struggle with complex problems, and evaluating responses to such problems can be just as challenging as answering them Wu et al. (2024a). Moreover, LLM-as-a-judge can introduce system biases, such as position bias, verbosity bias, and self-enhancement bias, which can undermine the fairness of the evaluation process Chen et al. (2024a); Zheng et al. (2023a).
3 WarriorCoder: Learning from Expert Battles
Figure 2: The diagram of learning from a battle between two chatbots.
In this section, we describe how WarriorCoder learns from expert battles. Different from the previous work, we build an arena for current expert code LLMs and organize competitions among them, other than expanding existing datasets through prompting the limited pool of proprietary LLMs. Each model takes advantage of what he has learned to challenge others under the evaluation of uninvolved judge models. The target model learns from the winner of these pairwise competitions and integrates the strengths of all the competitors.
3.1 Instruction Mining from Scratch
Considering a battle between A and B where A is on offense and B is on defense. The first step of our method is to use A’s strengths to challenge B, which makes it necessary to know what A has learned during its training process. However, nearly all the current open-sourced LLMs keep their core data proprietary. Inspired by Magpie Xu et al. (2024b), we design a completion-based method to mine the capabilities which the code LLMs have already mastered. Here we take Qwen2.5 Hui et al. (2024) as an example. A conversation about writing Python code in the chat template of Qwen2.5 is:
<|im_start|>system You are an AI assistant designed to provide helpful on Python coding problems.<|im_end|> <|im_start|>user
Write a Quicksort algorithm.<|im_end|>
<|im_start|>assistant
Here is the solution:
def quicksort(arr):
<|im_end|>
Such chat templates are pre-defined conversational structures to guide the interaction between the model and the user. Based on the LLMs’ strong completion abilities, we feed only the prefix of the chat template into them, prompting the LLM to generate the user queries:
<|im_start|>system You are an AI assistant designed to provide responses on Python coding problems.<|im_end|> <|im_start|>user
In this way, we can collect various instructions the model has already learned under various generation settings (different values of temperature and top-p). Unlike traditional synthetic data, our instructions are not synthesized by the models but directly sampled from their distributions, which avoids pattern overfitting and significant shifts in the output distribution Chen et al. (2024b). However, these instructions may be repetitive, ambiguous, unclear, or too easy. To address these concerns, we first deduplicate the data and adopt uninvolved LLMs in the arena to assess their difficulty. We divide the difficulty of instructions into four levels:
  • Excellent (9-10): For queries that are very clear, specific, and well-articulated. These queries are particularly challenging and excellently designed to assess the AI’s proficiency.
  • Good (6-8): For queries that are clear and specific instructions. These are not overly difficult to answer and moderately assess the AI’s capabilities.
  • Average (3-5): For queries that are fairly clear and specific instructions. These queries are easy to answer.
  • Poor (1-2): For queries that are ambiguous or unclear.
Only good and excellent instructions are considered during the following steps (the top half of Figure 2).
Then we compress the high-quality instructions for the efficiency of post-training. To ensure the diversity and representativeness of instructions, we employ KCenterGreedy algorithm Sener and Savarese (2018) to select the final instructions based on the embedding model - all-roberta-large-v1 Liu et al. (2019).
3.2 Win-Loss Decision
The defender is required to respond to the attacker’s question, while the attacker A must also provide an answer to its own query. Once both responses are collected, we use other LLMs in the arena to evaluate the correctness and helpfulness of the pairwise data (the bottom half of Figure 2). To ensure impartiality in the evaluation process, we establish a set of rules, including order shuffling, suspicion averting, and offensive-defense balance.
Order shuffling refers to the practice of randomizing the order in which responses from the attacker (A) and defender (B) appear in the evaluation prompt. This helps mitigate any positional bias that may arise if a particular position is favored by certain language models.
Suspicion averting ensures that competitors do not evaluate their own responses, preventing any potential bias in favor of their own generated answers.
Offensive-defense balance guarantees that all competitors have an equal number of offensive and defensive turns, maintaining fairness in the evaluation process.
Having the voting results, we can calculate the local score for each query $q$:
$x^{q}_{A>B}=\frac{t_{A}}{t_{A}+t_{B}}\quad x^{q}_{B>A}=\frac{t_{B}}{t_{A}+t_{B}}$ (1)
where $x^{q}_{A>B}$ and $x^{q}_{B>A}$ are A’s and B’s local score for $q$. $x^{q}_{A>B}$ represents the percentage of votes that candidate A receives, while $x^{q}_{B>A}$ similarly represents the percentage of votes that candidate B receives. $t_{A}$ and $t_{B}$ are the number of votes which A and B win.
However, relying solely on the local score to select the winner can be problematic. In some cases, a weaker model may receive more votes than a stronger one, even though its responses are not significantly better. This can occur because the local score may not fully capture the quality of the model’s performance, especially in situations where the voting is influenced by factors, such as randomness or bias from LLM judges.
To address this limitation, we propose considering both local contingency and global consistency in the decision-making process. Instead of directly basing our analysis on the immediate voting outcomes, we introduce the concept of the global score — specifically, the Elo rating Bai et al. (2022), which provides a more comprehensive reflection of a model’s relative performance over time and across various evaluations. The Elo rating system, originally developed to calculate the relative skill levels of players in two-player games (such as chess), has been successfully adapted to assess the performance of competitors in a range of competitive scenarios, including esports and other skill-based games.
By incorporating the Elo rating, we account for both local performance in individual contests and global performance across multiple rounds, providing a more robust and accurate measure of a model’s overall ability. This helps to mitigate the risk of weak models winning based on isolated, potentially unrepresentative votes:
$X^{Elo}_{A>B}=\frac{1}{1+{10}^{(R_{B}-R_{A})/400}}$ (2)
$X^{Elo}_{B>A}=\frac{1}{1+{10}^{(R_{A}-R_{B})/400}}$
where $X^{Elo}_{A>B}$ and $X^{Elo}_{B>A}$ indicate the expected probabilities of A defeating B and B defeating A, respectively. $R_{A}$ and $R_{B}$ are the Elo rating of A and B, which are updated dynamically and iteratively. Given the battle result of A and B on an instruction $q$, we update them by:
$R_{A}\leftarrow R_{A}+K\times(s^{q}_{A>B}-X^{Elo}_{A>B})$ (3)
$R_{B}\leftarrow R_{B}+K\times(s^{q}_{B>A}-X^{Elo}_{B>A})$
where $s^{q}_{A>B}$ and $s^{q}_{B>A}$ are the actual scores of player A and B (1 for a win, 0.5 for a draw, and 0 for a loss). The factor $K$ controls the sensitivity of rating changes. Here we set it to 0.005.
Based on Equation 1 and Equation 2, we can obtain the final score of A’s response for instruction $q$:
$e^{q}_{A}=\sum_{B\in Com\setminus A}\alpha X^{Elo}_{A>B}+(1-\alpha)x^{q}_{A>B}$ (4)
where $Com$ is the set of all the competitors and $\setminus$ is the subtraction operation. $\alpha$ is the coefficient to balance the local contingency and global consistency.
3.3 Final Training
Each item of the constructed dataset consists of an instruction, responses from various strong LLMs, and their corresponding scores, which supports multiple post-training methods like SFT and KTO Ethayarajh et al. (2024). It’s worth noting that we employ KTO instead of DPO Rafailov et al. (2023) because KTO utilizes all the responses while DPO only makes use of pairwised data. We select the response with the highest score as the gold output for SFT and set a threhold to distinguish the chosen and rejected completions for KTO.
4 Experiments
4.1 Experimental Details
Backbones
We use DeepSeekCoder-Base-6.7B Guo et al. (2024) to initialize the target model. As for the competitors of expert battles, we choose strong open-source LLMs including Llama-3.1 Dubey et al. (2024), Qwen2.5 Hui et al. (2024), and Mistral Jiang et al. (2023).
Datasets
To evaluate the effectiveness of our model, we conduct evaluations on HumanEval Chen et al. (2021), MBPP Austin et al. (2021), HumanEval+ Liu et al. (2023), and MBPP+ Liu et al. (2024). For a fair comparison, we use greedy decoding for generating answers to each question, consistent with previous work.
Experimental Settings
We conducted parallel training on 8 NVIDIA A800 80G GPUs for all backbone models. The global batch size was set to 512, and the number of epochs was set to 2. We used a learning rate of $1\times 10^{-5}$ and a weight decay of $3\times 10^{-7}$. Additionally, a WarmupLR scheduler with a warmup ratio of 0.2 was used. For both the offensive and defensive turns, the number was set to 70,000.
HumanEval HumanEval+ MBPP MBPP+ Rely on proprietary LLMs?
Proprietary Models Code-Davinci-002 47.0 - 58.1 - -
Code-Cushman-001 33.5 - 45.9 - -
GPT-3.5-Turbo 76.8 70.7 82.5 69.7 -
GPT-4-Turbo 90.2 86.6 85.7 73.3 -
Base Models DeepSeekCoder-Base (6.7B) 47.6 39.6 70.2 56.6 -
CodeLlama (7B) 37.8 35.4 59.5 46.8 -
StarCoder (15B) 34.1 29.3 55.1 46.1 -
Fine-tuned Models CodeT5+ (16B) 31.7 26.2 54.6 44.4
WizardCoder-CL (7B) 48.2 40.9 56.6 47.1
WizardCoder-SC (15B) 51.9 45.1 61.9 50.6
Magicoder-DS (6.7B) 66.5 60.4 75.4 61.9
MagicoderS-DS (6.7B) 76.8 70.7 75.7 64.4
Magicoder-CL (7B) 60.4 55.5 64.2 52.6
MagicoderS-CL (7B) 70.7 66.5 68.4 56.6
WaveCoder-DS (6.7B) 72.0 - 63.6 -
WaveCoder-CL (7B) 48.1 - 47.2 -
WaveCoder-SC (15B) 50.5 - 51.0 -
Ours WarriorCoder-DS (6.7B) 72.1 (+24.5) 66.0 (+26.4) 70.5 (+0.3) 57.8 (+1.2)
Table 1: The pass@1(%)on four benchmarks.
Baselines
The baselines consist of proprietary models, base models, and fine-tuned models.
Proprietary Models. These models, unlike open-source models, are developed, owned, and managed by a private entity or organization. They are trained on specialized or private datasets that are not publicly available to serve specific business needs or objectives. Access to these models is usually based on API calls. Proprietary Models include Code-Davinci-002, Code-Cushman-001, GPT-3.5-Turbo Ouyang et al. (2022) and GPT-4-Turbo OpenAI et al. (2024).
Base Models. They are the foundational, pre-trained models that serve as the core for further fine-tuning or adaptation to code tasks. Base Models include DeepSeekCoder-Base Guo et al. (2024), CodeLlama Rozière et al. (2023), and StarCoder Li et al. (2023).
Fine-tuned Models. These models are initially pre-trained on a large, general-purpose dataset and then fine-tuned on a smaller, code-specific dataset. This two-step process enhances the model’s performance on coding tasks by enabling it to leverage both broad general knowledge and more focused, domain-specific expertise. Fine-tuned models include CodeT5+, DeepSeek-Coder-Instruct Guo et al. (2024), WizardCoder Luo et al. (2024b), Magicoder Wei et al. (2024), and WaveCoder Yu et al. (2024). The suffixes -DS, -CL, and -SC denote the base models DeepSeekCoder-Base, CodeLlama-Python, and StarCoder, respectively.
4.2 Main Results
The overall results are summarized in Table 1. WarriorCoder achieves impressive performance, with a pass@1 rate of 72.1% (66.0%) on HumanEval (HumanEval+) and 70.5% (57.8%) on MBPP (MBPP+), surpassing nearly all other fine-tuned models. Notably, it shows a significant boost over the backbone model, particularly on HumanEval and HumanEval+ (with improvements of 24.5 and 26.4, respectively). WarriorCoder also outperforms both Magicoder-DS and WaveCoder-DS, which share the same backbone architecture and similar amounts of training data. This substantial performance gap highlights the effectiveness of our approach in generating higher-quality training data, providing a clear advantage over other models that rely on similar foundational setups.
Moreover, the performance of WaveCoder-DS on MBPP drops to 63.6% after fine-tuning, while our model improves to 70.5%. These results underscore WarriorCoder’s paradigm as a powerful data flywheel that effectively absorbs abilities from various expert code LLMs. Our method significantly enhances the capabilities of target models to generalize across the four benchmarks, showcasing its superiority in utilizing diverse data sources to drive performance improvements.
Moreover, previous data flywheels typically rely on augmentations and annotations generated using proprietary large language models (LLMs) and specially designed prompts. In contrast, our approach does not require human-generated prompts or proprietary LLMs. Despite this, we achieve competitive results that rival those of advanced code LLMs that depend on such resources. This underscores the quality of our data flywheel, demonstrating its capacity to significantly enhance the target model at a much lower cost.
4.3 Ablation Study
From HumanEval HumanEval+ MBPP MBPP+
Qwen 67.7 60.4 68.3 55.8
Mistral 70.1 64.0 67.7 54.8
All (ours) 72.1 66.0 70.5 57.8
Table 2: The results observed when learning from different teachers.
Table 2 presents the results observed when the target model learning from different teachers. Qwen and Mistral perform well on the HumanEval (HumanEval+) and MBPP (MBPP+) benchmarks, respectively. This indicates that each of these models enables the target model to learn a specific set of knowledge. However, WarriorCoder benefits from learning across all expert code LLMs, leveraging the strengths of each. As a result, it outperforms the others across all four benchmarks, demonstrating the advantages of integrating knowledge from multiple specialized teachers.
4.4 Data analysis
4.4.1 Diversity Analysis
Figure 3: The overlapping rate between the mined instructions and existing training datasets.
Figure 3 illustrates the overlap between the instructions mined from expert LLMs and those from existing training datasets, measured using the ROUGE score. The majority of the mined instructions exhibit a ROUGE score of less than 0.3, indicating that they are largely distinct from the instructions in these datasets. Notably, there are no mined instructions with a ROUGE score exceeding 0.6, further reinforcing the observation that the mined instructions are sampled from the internal distribution of expert LLMs but not simply replications or extensions of existing training data. As a result, these instructions demonstrate a higher degree of diversity, making them particularly valuable for training purposes, as they introduce novel examples that can enhance the capability of the target model.
4.4.2 Difficulty Analysis
Figure 4 illustrates the difficulty distribution of the mined instructions, providing insights into the experts’ internal knowledge. The majority of the instructions fall within the ’good’ level, with scores ranging from 6 to 8. Instructions rated at the ’excellent’ level (scores 9-10) represent only a small fraction of the total, suggesting that highly complex or advanced tasks are relatively rare. On the other hand, instructions with scores below 6 are excluded from the training set, as they tend to be either too easy or overly ambiguous. Such instructions are considered detrimental to the training stage, as they may not provide meaningful learning signals and could undermine the model’s performance and generalization ability.
Figure 4: The difficulty of mined instructions. As mentioned in Section 3.1, the difficulty of instructions is divided into four levels: excellent (9-10), good(6-8), average(3-5) and poor(1-2).
4.4.3 Case Study
Here we show two examples of mined instructions. The first has a score of 6:
I’m trying to combine two dictionaries into one and eliminate duplicate values for a given key the two dictionaries may have different structures, eg. dictA = {’cat1’:{’cat2’:’A’}, ’cat2’:{’cat3’:’B’}…} dictB = {’cat1’:{’cat2’:’C’}} Combining with a recursive function seems the best option but I can’t seem to get it right. **Please Help.**
And the second has a score of 9:
**Function to Get the Index of an Element in a 2D List (List of Lists)** ==================================== Create a function named ’get_index_2d’ that accepts three parameters: - a 2D list ’matrix’ - an element to search for ’target’ - a default value to return ’default’ If the ’target’ exists in the ’matrix’, the function should return its index (a tuple containing row, column). If the ’target’ is not found in the ’matrix’, the function should return the ’default’ value. """python def get_index_2d(matrix, target, default=None): # implement function """ ### Example Use Cases: | Matrix | Target | Default | Expected | | [[1,2],[3,4]] | 4 | None | (1,1) | | [[1,2],[3,4]] | 5 | None | None | | [[1,2],[3,4]] | 1 | (0,-1) | (0,0) | | [[’a’,’b’], [’c’,’d’]] | ’d’ | (-1,-1) | (1,1) |
Both instructions are clear and specific, providing a solid foundation for tackling the problem. However, the first instruction is more straightforward, with fewer requirements and a greater emphasis on practical tips and strategies to handle the task. In contrast, the second instruction introduces more complex conditions and does not include any guidance, making it more challenging to solve the problem.
5 Conclusion
This paper highlights the significant limitations of existing data flywheels for code models, which primarily rely on pre-existing datasets and annotations from a limited pool of proprietary LLMs. This dependency leads to a lack of data diversity and reinforces thr systemic biases. Even more concerning is the fact that many current open-source expert code LLMs keep their training data proprietary, further restricting access to diverse and high-quality data sources. To address these challenges, we propose WarriorCoder, which learns from the competitive knowledge of expert code LLMs, enabling the absorption of each expert’s strengths. Unlike existing methods that expand and refine datasets, our approach constructs training data from scratch, leveraging the expertise of the strongest models in a competitive setup. This allows us to achieve competitive performance without the need for pre-existing datasets and costly annotations. Furthermore, our approach is not limited to enhancing LLMs’ coding abilities and can potentially be applied to other complex tasks in the future.